OpenStudy (anonymous):

Log(x-3)/2 +log(x+2)/7 =0

OpenStudy (gorv):

u know property of log

OpenStudy (gorv):

\[\frac{ 1 }{ 2 }*\log(x-3)+\frac{ 1 }{ 7 }*\log(x+2)\]

OpenStudy (gorv):

\[a \log b= \log b^a\]